OpenStudy (anonymous):

y= 3 sin (2x-pi) find period and phase pellet

OpenStudy (anonymous):

phase pellet lol

OpenStudy (anonymous):

phase shift

OpenStudy (anonymous):

yeah i know

OpenStudy (anonymous):

period is \(\frac{2\pi}{b}\)which in your case is \(\frac{2\pi}{2}=\pi\)

OpenStudy (anonymous):

also you have shifted \(\pi\) units to the right

OpenStudy (anonymous):

actually maybe that last part is wrong since you can write it as \[3\sin(2(x-\frac{\pi}{2}))\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

easiest way to do it is to set \[2x-\pi=0\] and solve for \(x\)

OpenStudy (anonymous):

you get \(x=\frac{\pi}{2}\) in two steps